the velocity vector of a car has an x- component of 17m/sec and a y-component of -11m/sec. What are the magnitude and direction of the velocity vector?
1
Expert's answer
2020-10-25T18:29:33-0400

The velocity vector is:


"\mathbf{v} = (v_x,v_y) = (17,-11)\space m/s"

The magnitude is:


"v = \sqrt{v_x^2 + v_y^2 } = \sqrt{17^2 + (-11)^2 } \approx 20.25\space m/s"

The directon is:


"\theta = \arctan \left( \dfrac{v_y}{v_x} \right) \approx 327.1"

Answer. "v = 20.25\space m/s" and "\theta = 327.1" counting from the x-axis.